OpenStudy (anonymous):

Find the solution to the equation 64^(x – 3) = 4^2x.

OpenStudy (anonymous):

sorry 64 = 4^3

OpenStudy (anonymous):

4^3(x-3) = 4^2x 3x - 9 = 2x x = 9

OpenStudy (anonymous):

How'd you get 4^3(x-3) = 4^2x

OpenStudy (anonymous):

64 = 4^3